OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4). y = x + 8 y = x + 3 y = −x + 8 y = −x + 3

OpenStudy (anonymous):

they all have 4/3

OpenStudy (anonymous):

can u explain

OpenStudy (anonymous):

if two lines are parallel then they have the same slope. find the slope, and then use the order pair they gave you to write your new equation in standard form

OpenStudy (anonymous):

.-. eeeeek i dont know how to explain it lol and oh okay thanks robz8

OpenStudy (anonymous):

thanxs robz..

OpenStudy (mathstudent55):

Solve 4x − 3y = −12 for y. What do you get?

OpenStudy (anonymous):

-4?

OpenStudy (mathstudent55):

4x − 3y = −12 -4x -4x --------------- -3y = -4x - 12 Divide both sides by -3: y = (4/3)y + 4 Slope is 4/3, ok?

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

A parallel line to it must also have slope = 4/3. No choice given has a slope of 4/3. Are these the choices from another problem?

OpenStudy (nurali):

for parallel lines, they have the SAME slope, but different y-intercepts (the value of 'b' is different). 4x-3y= -12 -3y= -12 -4x y= 4 +4/3x In other words, your new equation is y= 4/3x +b So using y= 4/3x +b and the point (-3,4): 4= 4/3(-3/1) +B 4= -12/3 +b 4= -4 +b b= 8 Therefore y= 4/3x +8

OpenStudy (anonymous):

no i think nurali helped me thanks

OpenStudy (anonymous):

Parallel lines have same slopes therefore writing the given eq as below 4x − 3y = −12 i.e. 3y = 4x+12 i.e. y = 4/3 x+ 4 comparing it with y = mx + c, we get Slope of the required line m = 4/3 since req line passes through (−3, 4) i.e. x1=-3, y1= 4 Now the eq of line in slope point for is given as y-y1=m(x-x1) i.e. y-4=4/3(x+3) i.e. y=4/3x+8 and this is the required eq of line